pooja195 (pooja195):

A cheese burger has 200 fewer calories than a large fry. Two cheeseburgers and a large fry have 1100. How many calories are in each item?

OpenStudy (adamk):

y - 200 = cheese burger y = large fry 2(y - 200) + y = 1100

pooja195 (pooja195):

huh??

OpenStudy (adamk):

Solve for y in that equation. y - 200 will give you the amount of calories in a cheese burger and y will give you the amount of calories in a large fry.

pooja195 (pooja195):

x-y=200? 2x+1y=1100?

OpenStudy (adamk):

Nope. The amount of calories in the two foods aren't being subtracted from each other. It says "A cheese burger has 200 fewer calories than a large fry." So, y - 200 or x - 200. Whichever variable you want as the large fries. Then the large fries are just x or y. There is only one variable in this equation.

pooja195 (pooja195):

im so confued =/

OpenStudy (adamk):

Ok, say x = the calories in the large fry. The variable for calories in the cheeseburger will be x - 200. Our two variables: x = large fry x - 200 = cheese burger. Now, it says "Two cheeseburgers and a large fry have 1100." So, multiply 2 times the cheese burger calories plus the amount of calories in a large fry to get the total amount of calories. 2(x-200) + x = 1100

pooja195 (pooja195):

so this is a substitution problem?

OpenStudy (adamk):

I'm not sure.

pooja195 (pooja195):

2(x-200) + x = 1100 is the equation then?

pooja195 (pooja195):

is that the equation?

OpenStudy (adamk):

Yes

pooja195 (pooja195):

Thanks :)

OpenStudy (adamk):

Yep!